OpenStudy (anonymous):

If e^(rt) is given by e^(lambda*t)cosut+ie^(lambda*t)sinut. Show that d/dt (e^(rt)) =re^(rt) for any complex number, r If e^(rt) is given by e^(lambda*t)cosut+ie^(lambda*t)sinut. Show that d/dt (e^(rt)) =re^(rt) for any complex number, r @Mathematics

OpenStudy (jamesj):

e^(lambda*t)cosut+ie^(lambda*t)sinut \[ = e^{\lambda t + ui} \] \[ = e^{rt} \] Now, it's actually a slightly strange question this one because we should expect by general properties of the exponential function that \[ (d/dt) \ e^{rt} = r . e^{rt} \] But it looks like the question wants you to prove that this is the case here. So we have \[ e^{rt} = e^{\lambda t} (\cos ut + i \sin ut) \] Now differentiate and use the product rule.

OpenStudy (jamesj):

\[ (d/dt) (function \ above) \] \[ = \lambda e^{\lambda t} (\cos ut + i \sin ut) + e^{\lambda t} (-u \sin ut + ...) \] \[ = ... \] You take it from here.

OpenStudy (anonymous):

Why didn't you multiply e^rt into the brackets and just take the derivative?

OpenStudy (anonymous):

I got for the last bit e^λt (-usin(ut) + uicost (ut)) I still can't solve for it?

OpenStudy (jamesj):

Now you know that you want the entire expression to be equal to r e^(rt) = ( λ + iu) e^(rt) = ( λ + iu) e^λt (cos ut + i sin ut) So just manipulate the expression until you make it equal to that. It works, try it.

OpenStudy (anonymous):

How did you get = ( λ + iu) e^(rt) = ( λ + iu) e^λt (cos ut + i sin ut). You mean manipulate the expression above where I used the product rule right

OpenStudy (jamesj):

By definition e^rt = e^λt ( cos ut + i sin ut) Now you should also know that cos ut + i sin ut = e^iut hence e^rt = e^λt . e^iut = e^(λ+iu)t Thus r = (λ+iu) What you asked to prove using the property of real-valued functions and their derivatives, is that d/dt e^rt = r . e^rt Now, as we know what r is and e^rt is in terms of λ and u, we have that r . e^rt = (λ+iu) e^(λt (cos ut + i sin ut) ------- (*) Hence you need to show that d/dt e^rt = (*) So that's what we're doing. e^rt = e^λt ( cos ut +i sin ut) Differentiating that expression d/dt e^rt = .... stuff .... -------- (**) You want to show now that (**) = (*)

OpenStudy (anonymous):

Ohh I didn't realize cos ut + i sin ut = e^iut yeah I got it now. Thanks a lot for your help!

OpenStudy (anonymous):

Reviving this thread in case anyone else comes across it. I had this same problem. The key is to note that (-u sin(ut) + ui cos(ut)) = - i u (-i sin(ut) - cos(ut))

OpenStudy (anonymous):

That is, we can get away with multiplying that expression by 1 (obviously), and it just so happens that (-1)( i )( i ) = 1.